Find the area of a regular hexagon with the given measurement.
4-inch side
A = sq. in.

Answer by checkley77(12844)

A hexagon is a six sided figure made up of 6 equalateral triangles.
This equalateral triangle triangle has sides=4 inches.
To find the height we use the pythagorean formula.
a^2+b^2=c^2
2^2+b^2=4^2
4+b^2=16
b^2=16-4
b^2=12
b=sqrt12
b=3.4641 is the height.
Area=bh/2
area=4*3.4641/2
Area=13.8564/2
Area=6.9282 ans. for a single equalateral trianglr
6*6.9282=41.5692 in^2is the area of this hexagon.